# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ C.3 Gaussian Width bounds for Residual Networks (ResNets)
Note that the parameter vector $\mathbf{w}^{(l)}=\text{vec}(W^{(l)})$ and can be indexed with $j\in[m]$ and $j^{\prime}\in[d]$ when $l=1$ and $j^{\prime}\in[m]$ when $l\geq 2$. Then, we have
$\left[\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right]_{i,jj^{\prime}}$ $=\left[\frac{\partial\alpha^{(l)}}{\partial W^{(l)}}\right]_{i,jj^{\prime}}=\frac{1}{L\sqrt{m}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}~{}.$ (118)
For $l\in\{2,\dots,L\}$, noting that $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathbb{R}^{m\times m^{2}}$ and $\left\lVert V\right\rVert_{F}=\left\lVert\operatorname{vec}(V)\right\rVert_{2}$ for any matrix $V$, we have
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}$ $=\sup_{\|V\|_{F}=1}\frac{1}{L^{2}m}\sum_{i=1}^{m}\left(\phi^{\prime}(\tilde{\alpha}_{i}^{(l)})\sum^{m}_{j,j^{\prime}=1}\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}V_{jj^{\prime}}\right)^{2}$
$\leq\sup_{\|V\|_{F}=1}\frac{1}{L^{2}m}\|V\alpha^{(l-1)}\|_{2}^{2}$
$\leq\frac{1}{L^{2}m}\sup_{\|V\|_{F}=1}\|V\|_{2}^{2}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(a)}{\leq}\frac{1}{L^{2}m}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(b)}{\leq}\frac{1}{L^{2}m}\left(1+\frac{\beta}{L}\right)^{2(l-1)}$
@@
$\left\|\frac{\partial\alpha^{(1)}}{\partial\mathbf{w}^{(1)}}\right\|_{2}^{2}\leq\frac{1}{L^{2}m}\|\alpha^{(0)}\|_{2}^{2}=\frac{1}{L^{2}d}\left\lVert\mathbf{x}\right\rVert_{2}^{2}=\frac{1}{L^{2}m}~{},$
which satisfies the form for $l=1$. That completes the proof. ∎
We now restate and prove the main result for ResNets. See 6
@@
$f(\theta;x)=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m}}W^{(1)}\mathbf{x}))))~{},$ (120)
@@
$\alpha^{(l)}$ $=\alpha^{(l-1)}+\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (123)
@@
which is the output of the last layer or the so-called featurizer. Then, the Gaussian width of the average set is simply $w(\hat{A}_{n}^{(L)})$
For any hidden layer parameter $\mathbf{w}^{(l)}\in\mathbb{R}^{m^{2}}$ with $\mathbf{w}^{(l)}=\operatorname{vec}(W^{(l)})$, the gradient is in $\mathbb{R}^{m^{2}}$ and we have
@@
$=\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\left(\prod_{l^{\prime}=l+1}^{L}\frac{\partial\alpha^{(l^{\prime}}}{\partial\alpha^{(l^{\prime}-1)}}\right)\frac{\partial f}{\partial\alpha^{(l)}}~{}$ (126)
Define $\mathcal{Z}^{(l)}=\left\{Z^{(l)}\in\mathbb{R}^{m^{2}\times m}:\left\|Z^{(l)}\right\|_{2}\leq\frac{1}{L\sqrt{m}}\left(1+\frac{\beta}{L}\right)^{l-1}\right\}$, then according to Lemma 117, $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathcal{Z}^{(l)}$ with probability $1-\frac{2l}{m}$. Similarly, define =$\mathcal{B}=\left\{B\in\mathbb{R}^{m\times m}:\left\|B\right\|_{2}\leq 1+\frac{\beta}{L}\right\}$, then according to Lemma 114, $\frac{\partial\alpha^{(l^{\prime})}}{\partial\alpha^{(l^{\prime}-1)}}\in\mathcal{B}$ with probability $1-\frac{2}{m}$. Besides, define $\mathcal{V}$ the set of $\mathbf{v}$, then according to (95), $\mathbf{v}\in\mathbb{R}^{m}$ and $\left\|\mathbf{v}\right\|_{2}\leq 1+\rho_{1}$, so $w\left(\mathcal{V}\right)\leq c\sup_{\mathbf{v}\in\mathcal{V}}\left\|\mathbf{v}\right\|_{2}\sqrt{m}\leq c(1+\rho_{1})\sqrt{m}$.
@@
$w(\mathcal{H}^{(l)})$ $\overset{(a)}{\leq}C\sup_{Z^{(l)}\in\mathcal{Z}^{(l)}}\|Z^{(l)}\|_{2}\prod_{l^{\prime}=l+1}^{L}\sup_{B\in\mathcal{B}}\|B\|_{2}~{}w(\mathcal{V})$
$\overset{(b)}{\leq}C\frac{\mathbbm{1}}{L\sqrt{m}}\left(1+\frac{\beta}{L}\right)^{l-1}\left(1+\frac{\beta}{L}\right)^{L-l}c\left(1+\rho_{1}\right)\sqrt{m}$
$=C_{0}\left(1+\frac{\beta}{L}\right)^{L-1}$
$\overset{(c)}{\leq}C_{0}(1+\rho_{1})e^{\beta}~{}.$
where (a) follows from Lemma 9, (b) from the definition of $\mathcal{Z}^{(l)}$ and $\mathcal{B}$, and (c) from the fact that
$\left(1+\frac{\beta}{L}\right)^{L-1}\leq e^{\beta}~{}.$
Define ^$\hat{\mathcal{H}}_{n}^{(l)}=\left\{\frac{1}{n}\sum_{z\in z^{(n)}}\frac{\partial f}{\partial\mathbf{w}^{(l)}}(z)\mid W\in B_{\rho}^{\operatorname{Spec}}(\theta_{0})\right\}$, from a similar proof as (75) we have
$w(\hat{\mathcal{H}}_{n}^{(l)})\leq w(\mathcal{H}^{(l)})\leq C_{0}(1+\rho_{1})e^{\beta}$
Since
@@
According to Lemma 8, the Gaussian width of the average predictor gradient set is bounded by $\sum_{l=1}^{L}w(\hat{\mathcal{H}}_{n}^{(l)})+w(\hat{A}_{n}^{(L)})\leq C_{0}(1+\rho_{1})e^{\beta}+w(\hat{A}_{n}^{(L)})$, then there exists constant $c_{1},c_{2}$ such that the Gaussian width of the loss gradient set is bounded by $c_{1}w(\hat{A}_{n}^{(L)})+c_{2}(1+\rho_{1})e^{\beta}$. Therefore,
$w(\hat{\Xi}_{n}^{\text{rn}})\leq c_{1}w(\hat{A}_{n}^{(L)})+c_{2}(1+\rho_{1})e^{\beta}$
Since the spectral norm bound for each layer except the last layer holds with probability $(1-2/m)$ according to Lemma 4, and we use it $L$ times in the proof, the Gaussian width bound holds with probability $(1-2L/m)$. ∎
Appendix D Additional Experiment Results
Assumption 2 is a commonly used assumption [69, 90, 14], can be established in theory under standard assumptions, and we provide empirical evidence supporting it in Figure 4.
(a) ResNet18 on CIFAR-10
(b) FFN on CIFAR-10
(c) CNN on Fashion-MNIST
Figure 4: $L_{2}$ Norm of gradients for (a) ResNet18 on CIFAR-10, (b) FFN on CIAFR-10 and (c) CNN on Fashion-MNIST. Each experiment is repeated 5 times and we report the average and the maximum norms. The results demonstrate that the $L_{2}$ norm of gradients are bounded for these models.

### v2

@@ C.2 Gaussian Width bounds for Residual Networks (ResNets)
Note that the parameter vector $\mathbf{w}^{(l)}=\text{vec}(W^{(l)})$ and can be indexed with $j\in[m_{1}]$ and $j^{\prime}\in[d]$ when $l=1$, $j\in[m_{l}]$ and $j^{\prime}\in[m_{l-1}]$ when $l\geq 2$. Then, we have
$\left[\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right]_{i,jj^{\prime}}$ $=\left[\frac{\partial\alpha^{(l)}}{\partial W^{(l)}}\right]_{i,jj^{\prime}}=\frac{1}{L\sqrt{m_{l}}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}~{}.$ (105)
For $l\in\{2,\dots,L\}$, noting that $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathbb{R}^{m_{l}\times m_{l}m_{k-1}}$ and $\left\lVert V\right\rVert_{F}=\left\lVert\operatorname{vec}(V)\right\rVert_{2}$ for any matrix $V$, we have
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}$ $=\sup_{\|V\|_{F}=1}\frac{1}{L^{2}m_{l}}\sum_{i=1}^{m_{l}}\left(\phi^{\prime}(\tilde{\alpha}_{i}^{(l)})\sum_{j=1}^{m_{l}}\sum_{j^{\prime}=1}^{m_{l-1}}\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}V_{jj^{\prime}}\right)^{2}$
$\leq\sup_{\|V\|_{F}=1}\frac{1}{L^{2}m_{l}}\|V\alpha^{(l-1)}\|_{2}^{2}$
$\leq\frac{1}{L^{2}m_{l}}\sup_{\|V\|_{F}=1}\|V\|_{2}^{2}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(a)}{\leq}\frac{1}{L^{2}m_{l}}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(b)}{\leq}\frac{1}{L^{2}m_{l}}\prod_{k=1}^{l-1}\left(1+\frac{\beta_{k}}{L}\right)^{2}$
@@
$\left\|\frac{\partial\alpha^{(1)}}{\partial\mathbf{w}^{(1)}}\right\|_{2}^{2}\leq\frac{1}{L^{2}m_{1}}\|\alpha^{(0)}\|_{2}^{2}=\frac{1}{L^{2}m_{1}}\left\lVert\mathbf{x}\right\rVert_{2}^{2}=\frac{1}{L^{2}m_{1}}~{},$
which satisfies the form for $l=1$. That completes the proof. ∎We now restate and prove the main result for ResNets. See 5
@@
$f(\theta;x)=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m_{L}}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m_{1}}}W^{(1)}\mathbf{x}))))~{},$ (107)
@@
$\alpha^{(l)}$ $=\alpha^{(l-1)}+\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (110)
@@
By Lemma 10, we bound the Gaussian width of the overall gradient by the sum of the width of the gradients of the last layer parameters $\mathbf{v}\in\mathbb{R}^{m_{L}}$ and that of all the intermediate parameters $W^{(l)},\ell\in[L]$.
@@
which is the output of the last layer or the so-called featurizer. Then, the Gaussian width of the average set is simply $w(A^{(L)})$.
For any hidden layer parameter $\mathbf{w}^{(l)}\in\mathbb{R}^{m_{l}m_{l-1}}$ with $\mathbf{w}^{(l)}=\operatorname{vec}(W^{(l)})$, the gradient is in $\mathbb{R}^{m_{l}m_{l-1}}$ and we have
@@
Define $\mathcal{Z}^{(l)}=\left\{Z^{(l)}\in\mathbb{R}^{m_{l}\times m_{l}m_{l-1}}:\left\|Z^{(l)}\right\|_{2}\leq\frac{1}{L\sqrt{m_{l}}}\prod_{k=1}^{l-1}\left(1+\frac{\beta_{k}}{L}\right)\right\}$, then according to Lemma 104, $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathcal{Z}^{(l)}$ with probability $1-\sum_{k=1}^{l-1}\frac{2}{m_{k}}$. Similarly, define $\mathcal{B}^{(l)}=\left\{B^{(l)}\in\mathbb{R}^{m_{l}\times m_{l-1}}:\left\|B^{(l)}\right\|_{2}\leq 1+\frac{\beta_{l}}{L}\right\}$, then according to Lemma 101, $\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\in\mathcal{B}^{(l)}$ with probability $1-\frac{2}{m_{l}}$. Besides, define $\mathcal{V}$ the set of $\mathbf{v}$, then according to (82), $\mathbf{v}\in\mathbb{R}^{m_{L}}$ and $\left\|\mathbf{v}\right\|_{2}\leq 1+\rho_{1}$, so $w\left(\mathcal{V}\right)\leq c\sup_{\mathbf{v}\in\mathcal{V}}\left\|\mathbf{v}\right\|_{2}\sqrt{m_{L}}\leq c(1+\rho_{1})\sqrt{m_{L}}$.
@@
$w(\mathcal{H}^{(l)})$ $\overset{(a)}{\leq}C\sup_{Z^{(l)}\in\mathcal{Z}^{(l)}}\|Z^{(l)}\|_{2}\prod_{l^{\prime}=l+1}^{L}\sup_{B^{(l)}\in\mathcal{B}^{(l)}}\|B^{(l)}\|_{2}~{}w(\mathcal{V})$
$\overset{(b)}{\leq}C\frac{1}{L\sqrt{m_{l}}}\prod_{k=1}^{l-1}\left(1+\frac{\beta_{k}}{L}\right)\prod_{k=l+1}^{L}\left(1+\frac{\beta_{k}}{L}\right)\cdot c\left(1+\rho_{1}\right)\sqrt{m_{L}}$
=$C_{0}\frac{1+\rho_{1}}{L}\sqrt{\frac{m_{L}}{m_{l}}}\prod_{k\not=l}\left(1+\frac{\beta_{k}}{L}\right)~{}$
where (a) follows from Lemma 11, (b) from the definition of $\mathcal{Z}^{(l)}$ and $\mathcal{B}^{(l)}$. Since
@@
According to Lemma 10, the Gaussian width of the average predictor gradient set is bounded by
$\sum_{l=1}^{L}w(\mathcal{H}^{(l)})+w(A^{(L)})$ $\leq\sum_{l=1}^{L}C_{0}\frac{1+\rho_{1}}{L}\sqrt{\frac{m_{L}}{m_{l}}}\prod_{k\not=l}\left(1+\frac{\beta_{k}}{L}\right)+w(A^{(L)})$
$=C_{0}\frac{1+\rho_{1}}{L}\sqrt{m_{L}}\prod_{l=1}^{L}\left(1+\frac{\beta_{l}}{L}\right)\sum_{l=1}^{L}\frac{1}{\left(1+\frac{\beta_{l}}{L}\right)\sqrt{m_{l}}}+w(A^{(L)})$
then there exists constant $c_{1},c_{2}$ such that
$w(\Xi^{\text{rn}})\leq c_{1}w(A^{(L)})+c_{2}\frac{1+\rho_{1}}{L}\sqrt{m_{L}}\prod_{l=1}^{L}\left(1+\frac{\beta_{l}}{L}\right)\sum_{l=1}^{L}\frac{1}{\left(1+\frac{\beta_{l}}{L}\right)\sqrt{m_{l}}}$
Since the spectral norm bound for each layer except the last layer holds with probability $(1-\frac{2}{m_{l}})$ according to Lemma 6, and we use it $L$ times in the proof, the Gaussian width bound holds with probability $(1-\sum_{l=1}^{L}\frac{2}{m_{l}})$. ∎